undeadknight26 (undeadknight26):

last question!!! Visitors to a carnival can buy an unlimited-ride pass for $50 or an entrance-only pass for $20. In one day, 282 passes were sold for a total of $10,680. How many unlimited-ride passes were sold?

undeadknight26 (undeadknight26):

114 146 168 212***

OpenStudy (anonymous):

hmmmmmm..... w8 a moment

undeadknight26 (undeadknight26):

take your time please :D

OpenStudy (anonymous):

it is 168 exactly ... do you want proof?

undeadknight26 (undeadknight26):

not proof just want to know how you did it :D

OpenStudy (anonymous):

see you know that there were 282 tpasses sold right so the equation is x+y=282 right? where x and y are the respective types of tickets sold ...?? understood?

undeadknight26 (undeadknight26):

yes.

OpenStudy (anonymous):

now x+y=282 hence y=282-x

OpenStudy (anonymous):

got that?

OpenStudy (anonymous):

now see the amount collect is 10680 $ right that is =>50x+20y=10680 .... try to think a lil' here now substitite that equation which i gave you in the previos comment and viola ... you get x to be 168

OpenStudy (anonymous):

uhhmmmmmmmmmmmmmmmmm///// u ther?

undeadknight26 (undeadknight26):

Sorry my mom told me to get off the computer when i was done with the test...

OpenStudy (anonymous):

k lol.... bye dude!

OpenStudy (anonymous):

168